- Press Release
- August 15, 2022
NOAA/GOES SATOPS Morning Report 31 July 2000
SATOPS MORNING REPORT July 31, 2000
On Friday, July 28, at approximately 1510z (11:10am local) the Satellite
Operations Control Center (SOCC) experienced a power loss due to a GSA
technician cutting the power while working on a defected fire alarm. SOCC
operations failed over to the Command and Data Acquisition Stations.
No data was lost. At 1545z power to SOCC was restored and systems
returned to normal operations.
No change in the status of the GOES spacecraft.
On Saturday, July 29, GOES-10 frame breaks occurred at 0235z, 1038z
and 1948z due to loss of signal synchronization, and from 0435z to 0612z
and 1801z due to thunderstorms at Wallops CDA.
On Saturday, GOES-8 frame break occurred at 1817z due to east/west address
On Sunday, July 30, GOES-10 frame breaks occurred at 0227z due to equipment
configuration and at 0856z due to loss of signal synchronization at Wallops
CDA. GOES-10 frame breaks also occurred at 1717z and 1722z due to
thunderstorms at Wallops CDA.
No change in the status of the POES spacecraft.
On Friday, NOAA-12 rev 47811, L4A and L1A did not lock on frame sync.
L4A was 100% recovered but L1A was not recoverable due to limitations of
tape recorder management.
On Friday, NOAA-14 rev 28748, G1A was noisy from the spacecraft with
100% of the data recovered.
On Saturday, NOAA-15 rev 11489, G5B was noisy from the spacecraft with
97.8% of the data recovered.
On Saturday, NOAA-12 rev 47829, S5B and S1A did lock on frame
sync. S5B was 89.9% recovered and S1A was 100% recovered.
On Sunday, NOAA-14 rev 28775, G1B and G5A were noisy from the
spacecraft with 92% and 100% of the data recovered, respectively.
No change in the status of the DMSP spacecraft.
On Saturday, one OPS48 acquisition was reported as lost due to site
Questions or comments to firstname.lastname@example.org